Beam1a Tsys 19oct11
The galfa data
from 14oct11 was used to measure the ripple in the spectrum caused
by the fiber optics. On 19oct11 the fiber connectors were cleaned up
stairs and downstairs.
The galfacts data taken 19oct11 after the cleaning
showed that beam1a Tsys was 150
kelvins (.ps) (.pdf)
On 20oct11 the fibers for beam 1a were switched to:
- Black beam1a, red beam1b.
- The computation used: calValAvgK * calOff/(calon-caloff). I used
the cal value at 1450 MHz for the entire 172 MHz band centered at 1450
- The scalloping of beam1b is the Tsys variation with za. The dome
was moving 2 to 19.5 deg za.
After the fiber switch the Tsys for beam1a dropped backed to around 30
- beam1a cable -> beam6b fiber input
- beam6b cable -> beam1a fiber input (beam6b is not working)
- beam1a fiber receiver output -> beam6b cable
- beam6b fiber receiver output -> beam1a cable
Looking at the cal
deflection plots for a2130 from 20oct11 beam1a cal deflection now
has a +/- 2% ripple that varies with za.
- this was not present when fiber 1a was used.
- a2130 is using a winking cal. For the caldeflection to
change implies that the gain is changing (assuming tsys is
constant over the .2 second winking of the cal).
- At 2 and 20 deg za the turret floor is tilted by +/- 10 degs .I
wonder if there is something that is shifting with the tilt of the
floor causing the gain to change?